OpenStudy (anonymous):

A 25 ounce solution is 20 percent alcohol. If 50 ounces of water are added to it, what percent of the new solution is alcohol The book answer is 6 2/3 would someone explained how the book got this answer cause the way its explained in the answers threw me off

OpenStudy (anonymous):

You have 25oz of a solution of water and alcohol; 20% (or .2 in decimal) is alcohol so to figure out how much alcohol the solution is composed of you divide: 25oz\[\div\].2= 5oz of alcohol , from this we know that 20oz must be water to total to the solution amount of 25oz. 50oz of water is added to the 25oz solution which sums up to 75oz of solution (5oz of which is the alcohol). To figure out what percent the alcohol makes up in the solution you divide the amount of alcohol over the total amount of the solution: 5oz \[\div\] 75oz= 5/75= 1/15= .06666 which is 6.66% or 6 2/3 %

OpenStudy (anonymous):

ok that was pretty simple thankls